OpenStudy (anonymous):

Find an equation in standard form for the ellipse with the vertical major axis of length 16, and minor axis of length 10.

OpenStudy (anonymous):

is the center at the origin?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):



OpenStudy (anonymous):

so its x2/5+y2/8

OpenStudy (anonymous):

if that is the case, then it should be \[\frac{x^2}{5^2}+\frac{y^2}{8^2}=1\]

OpenStudy (anonymous):

yay!!

OpenStudy (anonymous):

denominators should be 25 and 64 respectively